OpenStudy (anonymous):

solve the following equation -17+5/3x=8; show work

OpenStudy (campbell_st):

add 17 to both sides 5/3x = 25 multiply both sides by 3x 5 = 75x divide by 75 x = 5/75 x = 1/15